OpenStudy (anonymous):

6 years ago, Eric's age was two less than four times his age 40 years ago. How old is Eric now?

OpenStudy (anonymous):

Eric's age is the unknown. Express his age 6 years ago relevant to his age 40 years ago.

OpenStudy (anonymous):

\[(x-6)+2=4(x-40)\]

OpenStudy (lgbasallote):

let x = eric's age x - 6 = eric's age 6 years ago x - 40 = eric's age 40 years ago "6 years ago, Eric's age was two less than four times his age 40 years ago" x - 6 = 4(x-40) - 2 x - 6 = 4x - 160 - 2 x - 6 = 4x - 162 162 - 6 = 4x - x 156 = 3x x = 156/3 x = 52

OpenStudy (lgbasallote):

does that help?

OpenStudy (anonymous):

yes ty

OpenStudy (anonymous):

u there?

OpenStudy (lgbasallote):

welcome :D

OpenStudy (anonymous):

can u help me with another one?

OpenStudy (lgbasallote):

ill try

OpenStudy (anonymous):

go to my new question

OpenStudy (anonymous):

k?

OpenStudy (lgbasallote):

where?